Agenda/Minutes on 2020/11/27
CRY+VIS Interface Meeting
Participants: Takahashi, Yano, Miyo, Ushiba, Ikeda, Hirata, Aso, Akutsu, Sato
- Leak test and removal of cryo-payload in EY.
- Evacuation of IYC. Height change: top chamber 1mm, BF 5mm. The jacks were touching the bottom of chamber.
- Jig for the FC painting. 2D drawings were ready.
- The connection/disconnection schedule.
- Disconnection: ITMY by 2/26
- Connection: ITMX after 3/2, ITMY after 3/30, ETMX after 1/12, ETMY after 5/4
- Acceptance check items.
PF actuator: Measurement of coefficients --> calculation of the noise (by the end of Jan.).
Participants: Takahashi, Yano, Miyo, Fabian, Terrence, Sato, Hirata, Aso, Washimi, Iguchi
- IYC chamber was evacuated. The height of top chamber went down by 1mm. BF LVDT by 5mm.
Week of the 23rd of November 2020
[Hirata, Washimi, Satou, Aso, Fabian]
On Tuesday morning we realized it was a lock nut that was too low the one causing the problem. See klog 15518.
- In the morning I moved the three nuts and the keystone moved down to about -1.2 mm.
- In the afternoon Hirata-san moved the nuts further up but the keystone didn't reach go beyond -1.2 mm.
- I checked the 3D-CAD and realized the OSEM flag and body impose a limit of ± 1.75 mm when the IM-IRM and OSEMs are perfectly aligned. In SR3 the limit is -1.470 mm.
- This sugguests -1.2 mm is the limit and it's determined by the overall alignment of the system. We should check where the limit is in SR2 and BS also.
- Health check of SRM payload is pending.
- The 30 mHz oscillation I talked about in last week's meeting is aliasing noise.
- I was looking at a slow channel that down-samples data from a fast channel, and the fast channel has high frequency noise (~250 Hz).
- This type of high frequency noise had already been noted by Type-A team and we had put low-pass filters in an inout fiter bank to avoid it. However, the filter had been turned off some time ago and the noise became obivous.
- A particular feature of the noise is that it's amplitude seems to depend on the position of the keystone.
In klog 15305 I showed the amplitude incresed when the keytone goes down. However, on Tuesday we saw the opposite. More analysis pending.
- When we began the installation of F0 fishing rod limit swtiches we realized some holes that were expected to be fully threaded were not threaded at the top. We quickly came up wit an additional plan and additional components have been ordered already. We'll finish installation next week.
We tested the polarity of thr F0 FR motor. We got the opposite I had reported in klog 15272. The reason is that in the script Ikeda-san had put a minus sign. It was agreed to remove the sign and change the polarity in hardware: we want the keystone to move up with a positive number of steps. This convention is alright, although the oppsite is common in motors we cannot access.
Satou-san realized he had made a mistake in the cabling diagram of the F0 yaw stepper motor (klog 15530). He quickly notified us and we will correct this error when we open the camber next week. (We did the one in SR3 correctly, klog 15441.)
- We removed 108 grams from the F0 keystone to compensate for the addiitonal mass of the limit switch components (~100 grams).
- Question for Ikeda-san and Aso-san: was it possible to test Ikeda-san's script to unjam F1 FR motor?
- Collateral damage: the fibre optic camera is not working. We tried to use it on Tuesday afternoon but it just keeps saying it's "Establishing connection".
We installed the limit switches in F0 FR and F0 yaw stepper motor and tested them using the medm. See klog 15441.
- We new F0 yaw stepper motor asembly is roughly 100 grams heavier, and the keystone doesn't have ballast mass on top to remove, so the load is heavier now. Later we will replace the top disc (1 kg) by a new lighter one.
- We had to remove some IP arch ballast masses to relocate a misplaced connector onto the IP table to allow us to connect the limit switches. We rearranged some cables also.
OpLev was installed in PRM.
- No progress
- Type-B remedy work is on going.
- Installation of the thermometers (pt100) onto the EQ stop frame.
- Procedure documents.
- Criteria in the acceptance test.
[Terrence] I would like to know the availability of the Type-B (or even Type-A) suspensions. I will be actively working on testing my methods on sensor corrections, complementary filters, and control optimization. I am also writing a paper on complementary filters so I will be collecting data from the suspensions.
The week from 11/30
- Takahashi: 11/30-12/4
- Hirata: 11/30-12/4
- Sato: 11/30-12/4
The week from 12/7
- Takahashi: 12/7-11
- Hirata: 12/7-11
- On 2020/12/4(Fri)